# Trust, Because You Can’t Verify: Privacy and Security Hurdles in Education Technology Acquisition Practices

### v1

@@ Abstract
The education technology (EdTech) landscape is expanding rapidly in higher education institutes (HEIs). This growth brings enormous complexity. Protecting the extensive data collected by these tools is crucial for HEIs. Privacy incidents of data breaches and misuses can have dire security and privacy consequences on the data subjects, particularly students, who are often compelled to use these tools. This urges an in-depth understanding of HEI and EdTech vendor dynamics, which is largely understudied.
To address this gap, we conduct a semi-structured interview study with $13$ participants who are in the EdTech leadership roles at seven HEIs. Our study uncovers the EdTech acquisition process in the HEI context, the consideration of security and privacy issues throughout that process, the pain points of HEI personnel in establishing adequate security and privacy protection mechanisms in service contracts, and their struggle in holding vendors accountable due to a lack of visibility into their system and power-asymmetry, among other reasons. We discuss certain observations about the status quo and conclude with recommendations to improve the situation.
@@ 1 Introduction
Digital technologies have become pervasive at higher education institutes (HEIs), integrated into every step of pedagogical and institutional processes[1, 2]. The huge amounts of data these education technologies (EdTech) collect, encompassing academic records, demographics, behavioral data, mobility and campus life, as well as financial, health, and employment records have made HEIs a lucrative target for cyber attacks: in 2023, data breaches at HEIs have cost an average of $3M$ USD [3]. Simultaneously, reports on the abuse of collected data by service providers and their affiliates for profiling, tracking, advertising, as well as direct selling of the data to data brokers, are on the rise [4, 5, 6]. Such practices have deep societal consequences, ranging from privacy violations to potential discrimination for employment and other benefits [7, 8, 1, 2]. In the long run, it can also create chilling effects from constant surveillance, undermining students’ freedom of expression [9, 10].
Prior research has uncovered privacy and safety issues of EdTech and concerns about those issues from students as well as instructors [11, 12, 13, 14, 15, 16], as they have no choice but to adopt institutionalized technologies [17, 18]. The lack of comprehensiveness in federal data protection laws (such as FERPA [19]) and their existing loopholes [20, 2] can be exploited by vendors to evade accountability [21, 22]. This motivates a need for in-depth scrutiny of the EdTech acquisition practices of HEIs and how they ensure security and privacy during procurement and throughout the life cycle of EdTech, which is largely unexplored. This paper addresses this gap through a semi-structured interview study ($n=13$) involving people who are in leadership positions at seven HEIs across US. Specifically, we aim to answer the following research questions:
@@
The only leverage that HEIs have over EdTech vendors during the acquisition process are law- and regulation-backed data policies (§ 5) such as FERPA [19] and HIPAA (Health Insurance Portability and Accountability Act [23]). While data breaches pose significant financial risks for HEIs, vendors often get away with only minor reputational hits (§ 5.1). Additional regulation, where it exists (e.g., at the state level), helps HEIs enforce additional security and privacy (and as a bonus, accessibility) requirements to EdTech vendors (§ 5.2). Though regulations help HEIs, ambiguities surrounding data breach liability limit HEIs’ ability to hold vendors accountable (§ 7.1).
@@
HEIs have almost no visibility into the security posture and practices of EdTech providers, resigning themselves to rely on questionnaire-based self-reported assessments (e.g., HECVAT [24]) to evaluate vendor’s risk profile and data management policies (§ 6.3). Some HEIs will be able to perform internal audits or penetration test, if the EdTech is directly integrated into the HEI system (§ 7.3). Notably, HEI visibility into sub-vendors is even worse, as HEIs lack authority over sub-vendors to undergo assessments, exposing them to potential security risks downstream (§ 6.4). Some small vendors and individual departmental EdTech acquisitions bypass institution-level assessments, leaving HEIs blind and powerless to how their data is used (§ 6.3).
@@
Contracts can only go so far in helping HEIs ensure data ownership and protection. Mining of data, especially as it relates to AI, is an increasing source of uncertainty and concern in HEI contracts with EdTech vendors. While contracts can limit vendor’s use of HEI data, the lack of visibility leave HEIs unaware of data misuse until after the fact (§ 7.1). EdTech’s off-boarding process is fraught with uncertainty, as HEIs lack the power and means to verify data deletion by vendors. Even contractual obligations often prove insufficient to ensure data erasure, leaving HEIs to rely on vendors’ promises and sometimes a confirmation message (§ 7.4).
We discuss the implications of these findings and conclude with suggestions to improve the status quo (§ 9).
@@ 2 Background and literature review
EdTech now manages nearly every aspect of academic activities [16, 1], with tech ecosystems constantly evolving with the availability of plug-and-play third-party ‘apps’ (such as app marketplaces for Zoom [^1] https://marketplace.zoom.us/ and Canvas [^2] https://app.learnplatform.com/marketplace), paving the way for additional data flows. Complexity brings vulnerabilities, as was seen when MOVEit, a popular file transfer tool, was compromised causing data breaches in almost 900 schools, exposing student’s personal information including name, IDs, date of birth, contact information, and Social Security number [26]. We lack an understanding of how HEIs conduct security assessments during procurement, how responsibilities for ensuring secure operations are distributed and maintained throughout the life cycle of EdTech, and how liability is negotiated in case of a privacy or security breach.
@@
Unfortunately, the US does not have any comprehensive data privacy law; the Family Educational Rights and Privacy Act (FERPA) is the main federal statute that governs the privacy of student data [2]. It requires HEIs to obtain consent before sharing “educational records” unless there is a “legitimate educational interest.” However, legal scholar Elana Zeide noted that institutions have almost complete authority to define what is “legitimate educational interest” and decide what data are protected by FERPA and what are not [20]. Russel et al. [2] noted that the revision to FERPA in 2008 gives more leverage to private companies by including them as “school officials” and allows disclosure of data to parties such as contractors, consultants, and volunteers, and once data enters the marketplace, it can be freely exchanged since FERPA does not apply to data brokers. Such loopholes allow vendor to use vague statements about what data can be collected, used, and for how long, as Paris et al. revealed after reviewing publicly available documents at Rutgers University [21]. For example, the contract with Canvas [^3] https://www.structure.com/canvas says that upon contract termination, the university will lose access to the data collected by Canvas, without mentioning if the data will be deleted or the vendor may continue to store and use it [21]. These findings are alarming, underscoring the need for an in-depth scrutiny of HEIs’ current EdTech acquisition practices—how they leverage federal and state legislature to guide vendor contracts and negotiations to protect data from breaches and misuse, and what challenges they face in doing so.
@@ 3 Methods
While our study received exempt status by our Institutional Review Board (IRB), we adhered rigorously to ethical and privacy standards recommended for human subjects studies. This included anonymizing personally identifiable information (PII) such as the participants and organizations names, and any other information that could reveal any identity during interview transcription.
@@
Participant Recruitment. We targeted EdTech leadership for their pivotal roles in guiding education technology adoption, implementation, and governance in HEIs. Specifically, University leadership provided operational insights, while information security officials offered expertise on security, privacy, and compliance matters.
Potential participants were identified through various US institutional websites and the authors’ professional contacts. The primary researcher contacted identified participants via a recruitment email that outlined study details, compensation, and instructions for interested parties. To expand the participant pool and include more individuals in leadership roles at HEIs, we also utilized snowball sampling [28], asking initial contacts to recommend additional participants. In total, we sent $51$ invites and interviewed $13$ participants from seven different US universities. Participant demographics are detailed in Table 1.
@@
Interview Design. Our interview questions explored three aspects of EdTech life cycle, including privacy policies and practices (regulatory compliance), the acquisition process (including criteria for selection and procurement processes), and challenges during post acquisition phase. We designed our questions to align with participants’ specific roles to gain focused and relevant insights. For example, while all participants discussed privacy policies, learning enterprises individuals provided detailed perspective on the EdTech acquisition process. Similarly, information security officials offered comprehensive insights into data management and security post acquisition. This approach enabled us to capture a comprehensive understanding of various perspectives and responsibilities within the EdTech ecosystem, offering a holistic view of how institutions balance educational goals with security and privacy standards. The interview questions for this study can be found in Appendix A.
@@
Pilot study. The research team collaboratively developed an initial interview questionnaire, which the primary researcher pre-tested through three pilot interviews. All interviewees were computer science (CS) faculty members experienced in EdTech use, including one who regularly conducts workshops and develops tutorials on the subject. They offered complementary perspectives, guiding adjustments in our approach and enhancing the quality of the data collected from the study participants [29]. These pilot interviews helped in clarifying questions, improving the language and interview flow (questions such as ‘factors that HEIs look for when acquiring EdTechs’ were expanded with more follow-up questions for greater depth), and revealing potential bias (such as maintaining a neutral tone). After multiple refinement rounds, we finalized the interview questions to align with our research goals.
@@
Data collection. The interviews followed a semi-structured approach, allowing the interviewer to skip or ask follow-up questions as needed, delve into deeper topics, and explore emerging themes. At the outset of each interview, participants provided verbal consent for their participation, with additional consent requested for audio recording purposes. The sequence and emphasis of questions varied based on a) the flow of the conversation and b) the participants’ role in EdTech acquisition. Participants were compensated \$$\$50$ through Amazon gift card for a 45-minute online conference call discussing security and privacy implications of EdTech acquisition. We recruited study participants between August 2023 and April 2024.
@@
Data Analysis. Anonymized transcripts underwent thematic analysis approach [30] using MAXQDA (a qualitative analysis software) [31]. To establish an initial codebook, two authors first independently analyzed four interviews, compared their initial codes, and collaboratively grouped them into broader categories. The authors used this initial codebook to independently analyze four new transcripts. They updated the codebook with new open codes and created a final codebook with $11$ overarching categories by refining already-created categories and resolved any disagreements [32]. These categories spanned EdTech life cycle at an HEI—from legal guidelines considered during acquisition to factors influencing the acquisition process to security assessments of technologies and post acquisition challenges. Authors used this final codebook to analyze three transcripts—two new ones and one of the previously coded transcripts—and achieved a inter-coder reliability score for Cohen’s Kappa ($\kappa>0.8$[33]. Since high agreement was reached, authors resolved any minor disagreements and continued analyzing the remaining transcripts independently. The final codebook can be found in Appendix B.
@@ 4 Overview of Results
Figure 1: EdTech acquisition and usage timeline learned from our interviews. Here, ● indicates a mandatory activity, and ○ indicates an optional one. Here we also present the major findings from our study.
Figure 1 illustrates the EdTech acquisition timeline provided by our participants, the key acquisition components and activities, as well as major findings pertaining to each phase in the EdTech acquisition process.
Guided by our research questions, we summarize the major findings here. To comprehend HEIs’ privacy posture concerning EdTech acquisition (RQ1, § 5), we scrutinize the legal frameworks (federal, § 5.1 state level, § 5.2, and international policies, § 5.3) and measures of creating governance models ( § 5.4) employed by HEIs. We uncover various challenges that HEIs encounter in enhancing their security and privacy posture, including the inadequacy of federal privacy laws, limited adoption of state laws, the complexities of navigating both intra- and international policies, and resource shortages for maintaining institutional governance policies.
We then examine how HEIs approach the EdTech acquisition process (RQ2) in Section 6. This includes examining a) the factors that are considered while acquiring an EdTech (§ 6.1), b) the importance of contracts to outline data ownership and security and privacy liabilities (§ 6.2), c) the existence of pre-acquisition security assessments and their challenges (§ 6.3), d) the additional challenges posed by the “web” of sub-vendors (§ 6.4), and e) the need for more collaborative groups where HEIs consult each other on various vendor experiences (§ 6.5).
Finally, we address challenges related to maintaining, auditing, and discontinuing EdTech (RQ3) post acquisition in Section 7). HEIs struggle determining liability (HEIs or vendors) in cases of vendor data breaches (§ 7.1). HEIs encounter challenges with auditability in vendor environments, relying on trust due to the opaque nature of these environments post-acquisition (§ 7.2). Participants highlighted difficulties with off-boarding EdTech, including issues with data access and deletion verification (§ 7.4). Nonetheless, HEIs continue to refine internal policies and processes to safeguard university data throughout the EdTech life cycle (§ 7.3).

### v2

@@ Abstract
The education technology (EdTech) landscape is expanding rapidly in higher education institutes (HEIs). This growth brings enormous complexity. Protecting the extensive data collected by these tools is crucial for HEIs as data breaches and misuses can have dire security and privacy consequences on the data subjects, particularly students, who are often compelled to use these tools. This urges an in-depth understanding of HEI and EdTech vendor dynamics, which is largely understudied.
To address this gap, we conducted a semi-structured interview study with $13$ participants who are in EdTech leadership roles at seven HEIs. Our study uncovers the EdTech acquisition process in the HEI context, the consideration of security and privacy issues throughout that process, the pain points of HEI personnel in establishing adequate protection mechanisms in service contracts, and their struggle in holding vendors accountable due to a lack of visibility into their system and power-asymmetry, among other reasons. We discuss certain observations about the status quo and conclude with recommendations for HEIs, researchers, and regulatory bodies to improve the situation.
@@ 1 Introduction
Digital technologies have become pervasive at higher education institutes (HEIs), integrated into every step of pedagogical and institutional processes[1, 2]. The huge amounts of data these education technologies (EdTech) collect, encompassing academic records, demographics, behavioral data, mobility and campus life, as well as financial, health, and employment records, have made HEIs lucrative targets for cyber attacks: in 2023, data breaches at HEIs have cost an average of $3M$ USD [3]. Simultaneously, reports on the abuse of collected data by service providers and their affiliates for profiling, tracking, advertising, as well as direct selling of the data to data brokers, are on the rise [4, 5, 6]. Such practices have deep societal consequences, ranging from privacy violations to potential discrimination for employment and other benefits [7, 8, 1, 2]. In the long run, it can also create chilling effects from constant surveillance, undermining students’ freedom of expression [9, 10].
Prior research has uncovered privacy and safety issues of EdTech and concerns about those issues from students as well as instructors [11, 12, 13, 14], as they have no choice but to adopt institutionalized technologies [15, 16]. The lack of comprehensiveness in federal data protection laws (such as FERPA [17]) and their existing loopholes [18, 2] can be exploited by vendors to evade accountability [19, 20]. This motivates a need for in-depth scrutiny of the EdTech acquisition practices of HEIs and how they ensure security and privacy during procurement and throughout the life cycle of EdTech, which is largely unexplored. This paper addresses this gap through a semi-structured interview study ($n=13$) involving people in leadership positions (including Chief Information Security Officers (CISOs), directors, and other senior individuals) at seven HEIs across the US. Specifically, we aim to answer the following research questions:
@@
The only leverage that HEIs have over EdTech vendors during the acquisition process is the law- and regulation-backed data policies (§ 5) such as FERPA (Family Educational Rights and Privacy Act [17]) and HIPAA (Health Insurance Portability and Accountability Act [21]). While data breaches pose significant financial risks for HEIs, vendors often get away with only minor reputational hits (§ 5.1). Additional regulation, where it exists (e.g., at the state level), helps HEIs enforce additional security and privacy (and as a bonus, accessibility) requirements to EdTech vendors (§ 5.2). Though regulations help HEIs, ambiguities surrounding data breach liability limit HEIs’ ability to hold vendors accountable (§ 7.1).
@@
HEIs have almost no visibility into the security posture and practices of EdTech providers, resigning themselves to rely on questionnaire-based self-reported assessments (e.g., HECVAT [22]) to evaluate vendor’s risk profile and data management policies (§ 6.3). Few HEIs perform internal audits or penetration tests for the HEI system when EdTech is directly integrated into the system (§ 7.3). Notably, visibility into sub-vendors is even worse, as HEIs lack authority over sub-vendors to undergo assessments, exposing them to potential security and privacy risks downstream (§ 6.4). Some small vendors and individual departmental EdTech acquisitions bypass institution-level assessments altogether, leaving HEIs blind and powerless to how their data is used (§ 6.3).
@@
Contracts can only go so far in helping HEIs ensure data ownership and protection. Data mining, especially relating to AI, is an increasing source of uncertainty and concern in HEI contracts with EdTech vendors. While contracts can limit vendors’ use of HEI data, the lack of visibility leaves HEIs unaware of data misuse until after the fact (§ 7.1). EdTech’s off-boarding process is fraught with uncertainty, as HEIs lack the power and means to verify data deletion by vendors. Thus, even when contracts require the vendors to delete institutional data after their service is discontinued, HIEs cannot verify whether data has been actually deleted; at best they can get a written confirmation (e.g., an email or other document) from the vendors stating that they have deleted all data (§ 7.4). We discuss the implications of these findings and conclude with suggestions to improve the status quo in Section 8.
Contributions.
Overall, our contributions are as follows:
  • We conduct the first qualitative study with HEI technology leadership to uncover the EdTech acquisition process and the underlying security and privacy posture in HEIs that dictates the process.
  • We provide insights into the nature of EdTech-related institutional policies of HEIs, including how these policies are enforced and monitored during EdTech acquisition and throughout their lifetime.
  • We highlight regulatory and technical limitations that hinder security and privacy compliance, especially in the post-EdTech acquisition phase.
@@ 2 Background and literature review
EdTech now manages nearly every aspect of academic activities [14, 1], with tech ecosystems constantly evolving with the availability of plug-and-play third-party ‘apps’ (such as app marketplaces for Zoom [^1] https://marketplace.zoom.us/ and Canvas [^2] https://app.learnplatform.com/marketplace), paving the way for additional data flows. Complexity brings vulnerabilities, as was seen when MOVEit, a popular file transfer tool, was compromised, causing data breaches in almost 900 schools. It exposed students’ data including name, IDs, date of birth, contact information, and Social Security number [24]. We lack an understanding of how HEIs conduct security assessments during procurement, how responsibilities for ensuring secure operations are distributed and maintained throughout the life cycle of EdTech, and how liability is negotiated in case of a privacy or security breach.
@@
Unfortunately, the US does not have any comprehensive data privacy law; the FERPA is the main federal statute that governs the privacy of student data [2]. It requires HEIs to obtain consent before sharing “educational records” unless there is a “legitimate educational interest.” However, legal scholar Elana Zeide noted that institutions have almost complete authority to define what is “legitimate educational interest” and decide what data are protected by FERPA and what are not [18]. Russel et al. noted that the revision to FERPA in 2008 gives more leverage to private companies by including them as “school officials” and allows disclosure of data to parties such as contractors, consultants, and volunteers; and once data enters the marketplace, it can be freely exchanged since FERPA does not apply to data brokers [2]. Such loopholes allow vendors to use vague statements about what data can be collected, used, and for how long, as Paris et al. reported after reviewing publicly available documents at Rutgers University [19]. For example, the contract with Canvas [^3] https://www.structure.com/canvas says that upon contract termination, the university will lose access to the data collected by Canvas, without mentioning if the data will be deleted or the vendor may continue to store and use it [19]. This situation also poses legal challenges for HEIs who enroll students from other countries with a stricter privacy policy. As reported in the Findings section, many HEIs we interviewed serve students from the European Union (EU) through online degree programs. Data about these students must be handled following the General Data Protection Regulation (GDPR [25]), which requires a lawful basis and stronger consenting requirements before sharing data with third parties and imposes restrictions on secondary data usage. Thus, there is a critical need for in-depth scrutiny of HEIs’ current EdTech acquisition practices—how they leverage federal and state legislature to guide vendor contracts and negotiations to protect data from breaches and misuse, and what challenges they face.
@@ 3 Methods
While our study received exempt status by our Institutional Review Board (IRB), we adhered rigorously to ethical and privacy standards recommended for human subjects studies. This included anonymizing personally identifiable information (PII) such as the participants’ and organizations’ names, and any other information that could reveal any identity during interview transcription.
@@
Participant recruitment. We recruited EdTech leadership for their pivotal roles in guiding education technology adoption, implementation, and governance at HEIs. Specifically, University leadership provided operational insights, while information security officials offered expertise on security, privacy, and compliance matters. Potential participants were identified through various US institutional websites and the authors’ professional contacts. The lead author contacted participants via a recruitment email that outlined study details, compensation, and instructions for interested parties.
We note the significant challenges associated with reaching this target population of leaders at HEIs with specific roles and professional experiences. To address these challenges, we adopted a flexible and iterative recruitment process. We conducted interviews with participants as they became available, continuously inviting new participants. This approach allowed us to engage with a diverse set of participants despite the difficulties in accessing the target population. To expand the participant pool, we also utilized snowball sampling [27], asking initial contacts to recommend additional participants.
In total, we sent 51 invites and interviewed 13 participants from seven US universities between August 2023 and April 2024. All our participants had over three years of experience working with EdTechs, with an average experience of $12$ years. Working with hundreds of EdTechs, our participants had first-hand experience dealing with vendors and associated challenges HEIs face in the EdTech landscape. Regarding educational background, four participants had degrees in business administration, and three had graduate degrees related to EdTech. Although we did not ask our participants directly about their educational background, we noted that information when reaching out to our participants from open sources. Information about the number of EdTechs handled was self-reported during the interview. More participant demographics are detailed in Table 1.
@@
Interview design. Our interview questions explored three aspects of the EdTech life cycle, including privacy policies and practices (regulatory compliance), the acquisition process (including criteria for selection and procurement processes), and post-acquisition challenges. We designed our questions to align with participants’ specific roles to gain focused and relevant insights. For example, while all participants discussed privacy policies, individuals from learning enterprises provided detailed perspectives on the EdTech acquisition process, while IT security officials offered comprehensive insights into data management and security post-acquisition. This approach enabled us to capture a comprehensive understanding of various perspectives and responsibilities within the EdTech ecosystem, offering a holistic view of how institutions balance educational goals with security and privacy standards. The interview questions for this study can be found in Appendix A.
@@
Pilot study. The research team collaboratively developed an initial interview questionnaire, which the primary researcher pre-tested through three pilot interviews. All interviewees were EdTech users, including one who regularly conducts workshops and develops tutorials on EdTech use. They offered complementary perspectives, guiding adjustments in our approach and enhancing the quality of the data collected from the study participants [28]. These pilot interviews helped clarify questions, improving the language and interview flow (questions such as ‘factors that HEIs look for when acquiring EdTechs’ were expanded with more follow-up questions for greater depth), and revealing potential bias (such as maintaining a neutral tone). After multiple refinement rounds, we finalized the questions and conducted the main study.
@@
Data collection. The interviews followed a semi-structured approach, allowing the interviewer to skip or ask follow-up questions as needed, delve into deeper topics, and explore emerging themes. At the outset of each interview, participants provided verbal consent for their participation, with additional consent requested for audio recording purposes. The sequence and emphasis of questions varied based on a) the flow of the conversation and b) the participants’ role in EdTech acquisition. During the interview, participants were provided sufficient time to ask clarifying questions and to gather their thoughts or articulate responses, ensuring their perspectives were accurately captured. To minimize interviewer bias, a neutral tone was maintained throughout the process, and the interviewer refrained from expressing personal opinions. Participants were compensated \$$\$50$ through an Amazon gift card for a 45-minute online conference call.
Interviews were conducted until data saturation was reached, signifying the point at which no new themes or insights emerge from the data [29], which is a standard in qualitative research. We collected and analyzed data in parallel and observed that saturation was achieved by the $13^{th}$ interview. This sample size is consistent with qualitative research guidelines, which suggest that saturation often occurs within 12-20 interviews [30]. More details about the analysis process are provided in the next section.
@@
Data analysis. Anonymized transcripts underwent a thematic analysis procedure [31] using MAXQDA (a qualitative analysis software) [32]. To establish an initial codebook, first, two authors independently analyzed four interviews, compared their initial codes, and collaboratively grouped them into broader categories. The authors used this initial codebook to independently analyze four new transcripts. They updated the codebook with new open codes and created a final codebook with $11$ overarching categories by refining already-created categories and resolving any disagreements through discussions [33]. These categories spanned EdTech life cycle at an HEI—from legal guidelines considered during acquisition to factors influencing the acquisition process to security assessments of technologies and post-acquisition challenges. Authors used this final codebook to analyze three transcripts—two new ones and one of the previously coded transcripts—and achieved an inter-coder reliability score for Cohen’s Kappa ($\kappa>0.8$[34]. A high inter-rater reliability score, such as this, indicates a consistent coding process, thereby supporting the credibility of the study’s findings [35] The authors resolved any minor disagreements and continued analyzing the remaining transcripts independently. The final codebook can be found in Appendix B.
@@
Figure 1: EdTech acquisition and usage timeline learned from our interviews. In light of our final codebook, we also summarize the important findings for each step, both for pre- and post-acquisition phases. Here, ● indicates a mandatory activity, and ○ indicates an optional one.
@@ 4 Overview of Results
Figure 1 illustrates the EdTech acquisition timeline provided by our participants, key acquisition components and activities, and major findings about each phase in the EdTech acquisition process. Section 5 details HEIs’ privacy posture (RQ1) by examining how federal ( § 5.1), state ( § 5.2), and international regulations ( § 5.3) inform institutional policies of creating governance models ( § 5.4). Section 6 examines EdTech acquisition approaches (RQ2, ), covering: a) factors considered in acquisitions (§ 6.1), b) the role of contracts for data ownership and liabilities (§ 6.2), c) pre-acquisition security assessments and their challenges (§ 6.3), d) issues with sub-vendors (§ 6.4), and e) the need for collaborative groups among HEIs to share vendor experiences (§ 6.5). Finally, challenges in maintaining, auditing, and discontinuing EdTech (RQ3) are discussed in Section 7. These include issues in determining liability in data breaches (§ 7.1), auditing vendor environments post-acquisition (§ 7.2), refining policies and internal auditing processes (§ 7.3. We note that while we included contextual information (e.g., mentioned pertinent regulations) and condensed data from several participants, all reported findings were solely derived from the interviews.